the sum of a number and it's square is 56. find the number.
Answer by checkley77(12844)

X+X^2=56
X^2+X-56=0
(X+8)(X-7)=0
X-7=0
X=7 ANSWER.
PROOF
7+7^2=56
7+49=56
56=56